OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x).

OpenStudy (anonymous):

Multiplicity refers to what it is raised to the power of. Roots/Solutions/Zeroes of -2 and 3 are (x+2)(x-3) Can you figure it out from this info?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

Roots = Solutions = Zeroes Thought you should know that these terms all apply to the same concept.

OpenStudy (anonymous):

Alright, so let's continue then.

OpenStudy (anonymous):

so those are already the solutions?

OpenStudy (anonymous):

Those are the factors. Ok so we need to know which factor the multiplicity of 2 is referring too. For example. The factor -2 is described by the (x+2) term. (It becomes zero when we plug in -2). If it's multiplicity is 2, that means it looks like (x-2)^2. Now if we are talking about the term 3, of which (x-3) describes the factor, where if we plug in 3 for x we get 0, right, if it is of multiplicity 2 then it looks like (x-3)^2.... is this making any sense?

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x). \[f(x) = (x-3)^{2}(x+2)\] OR \[f(x) = (x-3)(x+2)^{2}\] Depending on which term you are applying the multiplicity of 2 to.